OpenStudy (anonymous):

Differentiate the following function. f (x) = 7 x ^6 - 8 sin(x)

OpenStudy (anonymous):

f'(x) = 42x^5 - 8Cosx

OpenStudy (anonymous):

how did u get this? did you use the product rule

OpenStudy (anonymous):

because there is a minus sign inbetween the terms, you just differentiate the terms seperately, no need for the product rule :)

OpenStudy (anonymous):

if the problem had been: \[7x^{6}\sin(x)\] that calls for a product rule.

OpenStudy (anonymous):

yea good point joemath..

OpenStudy (anonymous):

is it product rule for this: \[8\sqrt{x}\sin x\]

OpenStudy (anonymous):

showing the work, your problem is this: \[\frac{d}{dx} (7x^{6}-8\sin(x)) = \frac{d}{dx}(7x^{6})-\frac{d}{dx}(8\sin(x)) = 7(6x^{5})-8(\cos(x))\] \[=42x^{5}-8\cos(x)\] Yes, it will be a product rule for that question.

OpenStudy (anonymous):

product rule is: \[(f(x)g(x))'=f'(x)g(x)+f(x)g'(x)\] \[f(x) = \sqrt{x} \Rightarrow f'(x) = \frac{1}{2\sqrt{x}}\] \[g(x)=\sin(x)\Rightarrow g'(x)=\cos(x)\] Plugging into the product rule we obtain: \[(\sqrt{x}\sin(x))'=\frac{1}{2\sqrt{x}}(\sin(x))+\sqrt{x}\cos(x)\] Multiply everything by 8 to get the correct answer :)

OpenStudy (anonymous):

Why do I multiply everything by 8....Why wasnt the 8 left in the equation

OpenStudy (anonymous):

Im using this rule: \[\frac{d}{dx}(cf(x)) = c(\frac{d}{dx}f(x))\] You can pull out constants when taking a derivative. So: \[(8\sqrt{x}\sin(x))' = 8(\sqrt{x}\sin(x))'\] I left out the 8, got an answer, and now i multiply by 8 make up for what i did earlier.

OpenStudy (anonymous):

so is it \[8\sin x+8\sqrt{x}\cos x\]

OpenStudy (anonymous):

sorry i forgot the 4/sqrt(x) in the front

OpenStudy (anonymous):

the answer I got leaving out the 8 was: \[(\sqrt{x}\sin(x))' = \frac{1}{2\sqrt{x}}\sin(x)+\sqrt{x}\cos(x)\] So putting the 8 back in would give: \[8(\sqrt{x}\sin(x))'=\frac{4}{\sqrt{x}}\sin(x)+8\sqrt{x}\cos(x)\] Yeah you got it :)

OpenStudy (anonymous):

g (t) = 6 t ^3 sin(t) for this one is the answer this: \[(6t^3)(\cos t)+(\sin t)(18t^2)\]

OpenStudy (anonymous):

thats correct :)

OpenStudy (anonymous):

its telling me its wrong tho

OpenStudy (anonymous):

hmm...i guess maybe it wants the answer simplified? there is some factoring we can do: \[6t^{3}\cos(t)+18t^{2}\sin(t) = 6t^{2}(tcos(t)+3\sin(t))\]

OpenStudy (anonymous):

You answer is correct though, dont think you did something wrong.

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

yup...it wanted it simplified

OpenStudy (anonymous):

thats dumb >.> at least you got it though :)

OpenStudy (anonymous):

ok one more! y = csc(θ) (θ + cot(θ))

OpenStudy (anonymous):

Lets see: \[f(\theta) = \csc(\theta) \Rightarrow f'(\theta)=-\csc(\theta)\cot(\theta)\] \[g(\theta)=\theta+\cot(\theta)\Rightarrow g'(\theta)=1-\csc^{2}(\theta)=cot^{2}(\theta)\] So putting this all in the product rule we have: \[(\csc(\theta)(\theta+\cot(\theta)))'=(-\csc(\theta)\cot(\theta))(\theta+\cot(\theta))+\csc(\theta)\cot^{2}(\theta)\]

OpenStudy (anonymous):

Lets simplify... one sec

OpenStudy (anonymous):

\[=-\theta \csc(\theta)\cot(\theta)-\csc(\theta)\cot^{2}(\theta)+\csc(\theta)\cot^{2}(\theta)=-\theta \csc(\theta)\cot(\theta)\]

OpenStudy (anonymous):

I hate trig functions =/

OpenStudy (anonymous):

i cant see the answer

OpenStudy (anonymous):

\[-\theta \csc(\theta)\cot(\theta)\]

OpenStudy (anonymous):

can you see it?

OpenStudy (anonymous):

yes i can thanks

OpenStudy (anonymous):

it say its wrong...this is frustrating

OpenStudy (anonymous):

ugh, let me look over my work to make sure i didnt do anything wrong first...

OpenStudy (anonymous):

I was a little hasty with my trig identity, uploading a pic of the solution in a sec

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

thanks alot it worked!

OpenStudy (anonymous):

Are you still there?